40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11634     Entry time: Tue Sep 22 16:42:39 2015
Author: ericq 
Type: Update 
Category: IOO 
Subject: Housekeeping 

I've moved the OAF MC2 signal path to go directly from c1oaf to c1mcs, so that the LSC being ON/OFF doesn't interfere with the MC length seismic feedforward. Since the FB is currently down, I can't do a full test, but looking at monitor points in StripTool indicates it's working as intended. 

I also cleaned up some LSC medm stuff; exposing the existing SRCL UGF servo, and removing a misleading arrow. This reminds me that I need to get calibration lockins back up and running...

ELOG V3.1.3-